Seth transport operates a fleet of Motor Lorries. The record for Lorry-4150 reveals the following information

Days Maintained 30

Days operated 25

Days Idle 5


Total hours operated 300

Total kms covered 2500

Total tonnage carried 200 (4 tonne – load per trip, return journey empty)

Total cost for the month 270000


Prepare a performance statement reflecting


a. Cost per day operated, cost per kilometer, cost per hour


b. Cost per round trip, Tonne Km, cost per commercial tonne –km.


Note formula plus calculations are required


1
Expert's answer
2021-05-20T04:58:02-0400

(a) cost per days operated 270000÷25 =10800

Cost per kilometer 270000÷2500= 108

Cost per hour 270000÷300= 900

(b) cost per trip 270000÷(200÷4)=5400